OpenStudy (oswaldmurphy):

Choose the equation below that represents the line passing through the point (−2, −3) with a slope of −6. y + 3 = −6(x + 2) y + 2 = 6(x + 3) y − 3 = −6(x − 2) y − 2 = 6(x − 3)

OpenStudy (oswaldmurphy):

@pooja195 @TheSmartOne

OpenStudy (ciarán95):

The general form of any linear equation is: \[y - y _{1} = m(x - x _{1})\] where m is our slope and (x1, y1) is some point which we know is on the line. In the question, we're given the slope and a point, so all you need to do is plug these into the above equation and you'll get the answer out! Hope that helps! :)

OpenStudy (oswaldmurphy):

So it is C?

OpenStudy (ciarán95):

Nope, but you're close! Just remember the negative signs in front of x1 and y1 in the general form of the equation....if we plug in -2 and -3, what will happen?

OpenStudy (oswaldmurphy):

It will be addition.

OpenStudy (oswaldmurphy):

So, A?

OpenStudy (ciarán95):

Yes, A is correct! Just to explain..... We can treat the negative signs in front of x1 and y1 as being a '-1'. So, we are multiplying our x1, which is -2, by -1 and we are also multiplying our y1, which is -3, by -1. \[y + (-1)(-2) = 6(x + (-1)(-3))\] MINUS TIMES MINUS IS PLUS, so our negative sign disappears and we get: \[y + 2 = 6(x + 3)\]